# Safe: Finding Sparse and Flat Minima to Improve Pruning

### v1

@@ 3.2 Augmented Lagrangian Based Approach
$x_{k+1},z_{k+1}$ $=\min_{x,z}\max_{\|\epsilon\|_{2}\leq\rho}f(x+\epsilon)+I_{\|\cdot\|_{0}\leq d}(z)$
$\hskip 60.00009pt+\frac{\lambda}{2}\|x-z+u\|_{2}^{2}$
$u_{k+1}$ $=\max_{u}\frac{\lambda}{2}\|x-z+u\|_{2}^{2}-\frac{\lambda}{2}\|u\|_{2}^{2}.$
@@ 3.6 Convergence Analysis
Table 1: Perplexities (mean±$\pm$±std) of LLaMa models pruned to various sparsity levels using different methods. Safe archives competitive performance, while ${}^{+}_{\text{}}$ outperforms baselines across all settings.
@@ 4.5 Comparison with Other SAM-based pruner
IMP+SAM (Na et al., 2022) involves applying SAM during iterative magnitude pruning (Liu et al., 2024; Frankle & Carbin, 2019). While it was initially introduced for language model finetuning, we adapt this method to image classification and use the same training epochs and sharpness-minimization hyperparameters as Safe to ensure fair comparison. Pruning is performed every 10 epochs with sparsity increasing either linearly, following Na et al. (2022), or cubically (Zhu & Gupta, 2017), with the latter yielding better performance.
[…]
Table 4: Comparison with IMP+SAM and CrAM on ResNet-20/CIFAR-10. Safe outperforms over all sparsity levels.
@@
CrAM 93.25±0.08 90.18±1.80 69.53±12.36 45.17±20.86 10.00±0.00
CrAM+ 94.21±0.19 93.62±0.06 91.75±0.41 88.82±0.18 81.30±0.56
[…]
Safe 93.44±0.01 92.59±0.09 89.58±0.10 87.47±0.07 79.55±0.13
Safe ${}_{\text{Multi}}$ 92.74±0.09 91.98±0.13 88.89±0.13 86.42±0.10 83.00±0.15
[…]
@@ Appendix C Detailed Results for Image Classification Tasks
Appendix D Additional Comparison with IMP+SAM on LLM pruning
@@ Appendix E Computation Cost Analysis LLM Pruning
E.1 Theoretical Computation Complexity
We compare the time complexity in terms of hidden dimensions $d$, number of data $N$ or batch size $b$, number of iterations $k$, and number of layers in a single transformer block $L_{B}$ to observe how the computation cost scales. The results are given in Table 9.
@@ F.2 Effects of Batch-norm Tuning
F.3 Effects of dual-update interval $K$
(a) Dense validation accuracy
(b) Sparse validation accuracy
(c) Distance to constraint
F.4 Effect of $\lambda$ Scheduling

### v2

@@ 3.2 Augmented Lagrangian Based Approach
$x_{k+1},z_{k+1}$ $=\underset{x,z}{\operatorname{argmin}}\max_{\|\epsilon\|_{2}\leq\rho}f(x+\epsilon)+I_{\|\cdot\|_{0}\leq d}(z)$
$\hskip 50.00008pt+\frac{\lambda}{2}\|x-z+u_{k}\|_{2}^{2}$
[…]
@@ 3.6 Convergence Analysis
Table 1: Perplexities (mean±$\pm$±std) of LLaMa models pruned to various sparsity levels using different methods. Safe achieves competitive performance, while ${}^{+}_{\text{}}$ outperforms baselines across all settings.
@@ 4.5 Comparison with Other SAM-based pruner
IMP+SAM (Na et al., 2022) involves applying SAM during iterative magnitude pruning (Liu et al., 2024; Frankle & Carbin, 2019). While it was initially introduced for language model finetuning, we adapt this method to image classification and use the same training epochs and sharpness-minimization hyperparameter search range as Safe to ensure fair comparison. Pruning is performed every 10 epochs with sparsity increasing either linearly, following Na et al. (2022), or cubically (Zhu & Gupta, 2017), with the latter yielding better performance.
Compression-Aware Minimizer (CrAM) (Peste et al., 2022), on the other hand, extends upon the robust optimization principles of SAM to train compressible models by encouraging the models to maintain strong post-compression performance under the presence of small perturbations as $\min_{x}\max_{\|\epsilon\|\leq\rho}\,f(C(x+\epsilon))$ given some compression operation $C$. This leads to the CrAM update rule $x_{t+1}=x_{t}-\eta\nabla f(C(x+\rho\nabla f(x)))$. Along with this, we additionally compare with CrAM+, a variant introduced by Peste et al. (2022) that simply adds the original gradient $\nabla f(x)$ in the update as $x_{t+1}=x_{t}-\eta\left[\nabla f(C(x+\rho\nabla f(x)))+\nabla f(x))\right]$. We run these using the optimal hyperparameters as suggested in Peste et al. (2022). It should be noted that the additional technique of CrAM+ is somewhat auxiliary to the core robust optimization mechanism that connects CrAM to Safe, and thus, care must be taken when associating their performance gains with the central strategy that defines CrAM. For better comparison, we extend this strategy to Safe by adding the gradient computed at projected point $\nabla f(C(x))$ during iterative $x$-minimization of Safe, which we denote as Safe${}_{\text{+SG}}$.
Table 4: Comparison with IMP+SAM, CrAM, and CrAM+ on ResNet-20/CIFAR-10. Safe${}_{\text{+SG}}$, which extends Safe using a similar technique as CrAM+, outperforms most baselines at moderate sparsity and all baselines at extreme sparsity.
@@
CrAM 90.18±1.80 69.53±12.36 45.17±20.86 10.00±0.00
CrAM+ 93.62±0.06 91.75±0.41 88.82±0.18 81.30±0.56
Safe 92.59±0.09 89.58±0.10 87.47±0.07 79.55±0.13
Safe${}_{\text{+SG}}$ 92.40±0.06 90.09±0.13 89.13±0.06 85.85±0.09
As shown in Table 4, Safe and Safe${}_{\text{+SG}}$ outperforms IMP+SAM and CrAM, where Safe${}_{\text{+SG}}$ demonstrates competitive performance to CrAM+ on moderate sparsity and outperforms it in extreme sparsity. We suspect that the pruning operation in IMP+SAM may be overly abrupt, potentially hindering the benefits of sharpness minimization. We additionally observe similarly strong performance over IMP+SAM on language model pruning in Section D.2. Conversely, although CrAM+ achieves competitive results, its benefits fails to extend to extreme sparsity. More crucially, while Safe achieves reliable performance without much additional techniques, CrAM, by itself, performs poorly in all sparsities, depending heavily on auxiliary techniques to drastically improve performance. This suggests that caution is warranted when attributing these gains to the effectiveness of their robust optimization formulation inspired by sharpness-minimization, which is more likely provided through the use of both the original dense gradient and the sparse gradient computed at the projected point from CrAM+. A similar trend is observed in Safe${}_{\text{+SG}}$, further supporting this interpretation. Safe, on the other hand, delivers competitive performance without relying on these additional techniques, highlighting the intrinsic effectiveness of its smooth penalization via the augmented Lagrangian and split-variable structure of the ADMM framework to jointly balance sharpness minimization and sparsity. We provide additional comparison with other variants of CrAM in Section D.1.
@@ Appendix C Detailed Results for Image Classification Tasks
Appendix D Additional Comparison with SAM-based pruners
Table 8: Comparison with other variants of CrAM on ResNet-20/CIFAR-10, where we also apply similar techniques to Safe.
Sparsity
Method 95% 98% 99% 99.5%
CrAM 90.18±1.80 69.53±12.36 45.17±20.86 10.00±0.00
CrAM+ 93.62±0.06 91.75±0.41 88.82±0.18 81.30±0.56
[…]
Safe 92.59±0.09 89.58±0.10 87.47±0.07 79.55±0.13
Safe${}_{\text{+SG}}$ 92.40±0.06 90.09±0.13 89.13±0.06 85.85±0.09
[…]
Safe${}_{\text{+SG,Multi}}$ 92.17±0.12 90.79±0.04 88.11±0.14 84.92±0.14
@@ Appendix D Additional Comparison with SAM-based pruners
D.1 Other variants of CrAM
[…]
D.2 IMP+SAM on Language model pruning
@@ Appendix E Computation Cost Analysis LLM Pruning
E.1 Theoretical Time Complexity
We compare the time complexity for pruning a single transformer block with Safe, SparseGPT, Wanda, and ALPS in terms of hidden dimensions $d$, number of data $N$ or batch size $b$, number of iterations $k$, and number of layers in a single block $L_{B}$ to observe how the computation cost scales. The results are given in Table 11.
E.2 Wall-clock Time
@@ F.2 Effects of Batch-norm Tuning
F.3 Effect of $\lambda$ Scheduling
@@
F.4 Effects of dual-update interval $K$
(a) Dense validation accuracy
(b) Sparse validation accuracy
(c) Distance to constraint